Graph the inequality
Graph the inequality
Solve the compound inequality
Solve the compound inequality
Solve the absolute value equation
Solve the absolute value equation
Graph the following inequality
Solve the absolute value expression
Answer #1
QuestionDetails:
Graph the inequality
Graph the inequality
Solve the compound inequality
this is a number line THIS IS ALLREAL NUMBERS if "and" not "or" then between 4 and 10
Solve the compound inequality this means that youranswer (usually a range) must make both true at the sametime
IS THIS TYPED CORRECTLY? this is x>= 1 and x >=5 to meet both
x>=5 [this is usually with opposite inequalities, checkthe problem]
Solve the absolute value equation
get the absolute valueisolated   |9n + 1 | = -2
since all absolute values are 0 or positive...there is nosolution
Solve the absolute value equation
think what value for x will make6-9x = 23 OR   6 - 9x = -23 the two answers are thesolution (note these answers may be negative, positive, or0
Graph the following inequality  GRAPHING INEQUALITIES:
1. FIND POINTS AS IF THE INEQUALITY WAS AN EQUATION
2. determine if points on the line are included (if (< or=) or (>or =) use solid line OR not included (noequal sign with inequality) use dotted line.
3. test each side of the line (using x- or y- intercepts isoften quick). test both side as a check...
Shade the side of the line where the point wastrue
same as x + y > 1 graph with dottedline y = -x +1, test points and shade true side
Solve the absolute value expression
same as above |x - 9| = 1 sowhat values for x will make (x - 9) = 1 or (x - 9 ) = -1
that's it.
This is 10 separate questions not one, so please rategenerously